OpenStudy (anonymous):

Let f(x)=x^2 +2x-15 . Determine the values of x for which a) f(x)=-12

jimthompson5910 (jim_thompson5910):

f(x)=x^2 +2x-15 -12=x^2 +2x-15 0 = x^2 +2x-15 +12 0 = x^2 +2x-3 x^2 +2x-3 = 0 I'll let you finish

OpenStudy (anonymous):

THe answer says its a x,y coordinate

jimthompson5910 (jim_thompson5910):

Solve for x. Once you get the solutions for x, you can find the corresponding values for y.

OpenStudy (anonymous):

I am sorry, I am just very confused.

OpenStudy (anonymous):

COuld you help me any further or is that all?

OpenStudy (anonymous):

I don't get how it fits into a co ordinate

OpenStudy (anonymous):

try to solve that quadratic for x (factor or use the quad formula)

OpenStudy (anonymous):

would quadratic be easier?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

o ohkay thank you

OpenStudy (anonymous):

i Haven't factored in forever how would you factor: x^2+2x-3?

OpenStudy (anonymous):

what are the factors of -3?

OpenStudy (anonymous):

why does that matter

OpenStudy (anonymous):

factored form of a quadratic looks like (x +a)(x +b)

OpenStudy (anonymous):

Yes I understand that part

OpenStudy (anonymous):

notice that a times b is going to have to equal the constant term in your quadratic

OpenStudy (anonymous):

your quadratic's constant term is -3

OpenStudy (anonymous):

So those two have to equal that?

OpenStudy (anonymous):

so if you can find the factors of -3 you'll have a few guesses as to what a and b should be, and you can check to see if your guesses work or not

OpenStudy (anonymous):

1, -3

OpenStudy (anonymous):

how would i write the part of (x + a) (x+b)

OpenStudy (anonymous):

(x+1) (x-3)?

OpenStudy (anonymous):

(x +3)(x -1) =?

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

Oh I see and the -3 constant can just disapear?

OpenStudy (anonymous):

no, of course not.

OpenStudy (anonymous):

expand that factored form and see if it matches the equation you are trying to factor. If it does, you have found the factors of the equation.

OpenStudy (anonymous):

if it doesn't then try the factors: (x -3)(x +1) =?

OpenStudy (anonymous):

how would you expand once more?

OpenStudy (anonymous):

because the answer is incorrect it should be the opposite signs on the numbers

OpenStudy (anonymous):

?

OpenStudy (anonymous):

expand like this: (x +3)(x -1) = x(x-1) + 3(x-1)

OpenStudy (anonymous):

finish that up and see if it matches the original equation you're trying to factor.

OpenStudy (anonymous):

i got 5x-2

OpenStudy (anonymous):

that is not the orignal equation

OpenStudy (anonymous):

x*x =? x*(-1) =? 3*x =? 3*(-1) =?

OpenStudy (anonymous):

when you get done, add any like terms.

OpenStudy (anonymous):

Its 5x -2

OpenStudy (anonymous):

try again

OpenStudy (anonymous):

oh my god

OpenStudy (anonymous):

4x-3?

OpenStudy (anonymous):

what is x times x?

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

what about the imaginary 1 infront

OpenStudy (anonymous):

wait nvm i'm dumb

OpenStudy (anonymous):

3*3 = 2*3? 4*4 = 2*4? 5*5 = 2*5?

OpenStudy (anonymous):

3*3 = 3^2 4*4 = 4^2 5*5 = 5^2

OpenStudy (anonymous):

*= to multiplying?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

use * or ( ) for multiplication from now on or you'll get multiplication confused with an 'x'

OpenStudy (anonymous):

how is 3*3= 2*3 as you wrote above

OpenStudy (anonymous):

3x3 is 9 and 2x3 is 6

OpenStudy (anonymous):

it's not. That's what you were saying. x*x isn't 2*x I was showing you how to prove it to yourself by using actual numbers. 3*3 is clearly not 2*3 it's 3^2

OpenStudy (anonymous):

okay well anyways back to the question

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

let me know when you've found x*x =? x*(-1) =? 3*x =? 3*(-1) =? and added the like terms

OpenStudy (anonymous):

x^2+2x-3?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

it matches the equation you were trying to factor. So therefore you have found the factored form.

OpenStudy (anonymous):

(x-3)(x+1) is the factored form of x^2+2x-3 Now we use it to find (x-3)(x+1) =0

OpenStudy (anonymous):

what values of x make that last statement true?

OpenStudy (anonymous):

3 and -1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so those are the two solutions to x^2 +2x-3 = 0

OpenStudy (anonymous):

Okay but in my book it says the answers are -3, and 1

OpenStudy (anonymous):

thats why i'm wondering

OpenStudy (anonymous):

Yeah we were actually expanding (x+3)(x-1) not (x-3)(x+1) so (x+3)(x-1) =0 that's the one you need to use to find x

OpenStudy (anonymous):

-3 and 1

OpenStudy (anonymous):

So what is the distinguishing factor for me to know if i should use -3 and 1 or 3 and -1

OpenStudy (anonymous):

just the steps we went over... try expanding (x-3)(x+1) and see what you get. You won't get x^2 +2x-3

OpenStudy (anonymous):

after you do a few you won't often need to go through these steps. It'll just be something you can do in your head by looking at the problem for a minute or two.

OpenStudy (anonymous):

so you just knew that it was -3 and 1 without doing none of those steps :/

OpenStudy (anonymous):

well, here's a little trick or short cut: x*x =? x*(-1) =? 3*x =? 3*(-1) =? look at that part again... see how there are two terms with x in them? (-x and 3x) those have to add to give you the 'x' term in your original...

OpenStudy (anonymous):

so your two terms a and b in (x+a)(x+b) have to multiply to give you -3 and add to give you 2

OpenStudy (anonymous):

I Just did (x+1) (x-3) and the oppostie way with a formula called FOIL (First, outside, inside, last) and both gave me x^2+2x-3

OpenStudy (anonymous):

bit of practice and you can do that part pretty quickly in your head, if you can also find the factors of the constant term in your head eg.: x^2 -6x +9 a*b = 9 a+b =-6 factors of 9 we will use that add to -6 must be -3 and -3 so (x-3)(x-3) is the factored form and the solutions are x= 3,3

OpenStudy (anonymous):

But yes I understand now that they must multiply to give you the constant and add to give you the number infront of the x alone

OpenStudy (anonymous):

yep:) exactly. and hopefully you also know why it has to work like that now:)

OpenStudy (anonymous):

Yes, so it came down to factoring :), and this would have worked with the quadratic formual as well?

OpenStudy (anonymous):

of course. lot more work however...

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

I'm just wondering are you a teacher or just someone well educated on the topic?

OpenStudy (anonymous):

I work for a math and physics education company. (no, not Kahn Academy)

OpenStudy (anonymous):

Oh I see, and you enjoy to help students?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Okay, well thank you very much again, I appreciate what you're doing keep it up :)

OpenStudy (anonymous):

No problem!